how do yew do this promblem
it is the linear system
please help me
here is the promblem
4x+y=12
y=-8

Answer by rapaljer(4551)

4x+y=12
y=-8

In this case, just substitute the y=-8 into the first equation:
4x+(-8) = 12
4x-8=12
4x=12+8
4x=20
x=5

Therefore the solution is (5,-8).

R^2